OpenStudy (anonymous):

Simplify completely x^2+4x-45/x^2+10x+9 and find the restrictions on the variable.

OpenStudy (anonymous):

x^2+4x-45/x^2+10x+9 =(x+9)(x-5)/(x+9)(x+1)=(x-5)/(x+1)

OpenStudy (anonymous):

thanks